# LeanAgent: Lifelong Learning for Formal Theorem Proving

### v2

@@ A.6 Curriculum Learning Analysis
Curriculum Learning. Prior work suggests that curriculum learning guides the learner towards better local minima in non-convex optimization problems [6]. Crucially, theorem proving involves navigating a highly non-convex optimization landscape, especially for complex mathematical statements. The space of possible proofs is vast and complex, with many potential dead ends and suboptimal solutions to parts of a proof. This makes theorem proving an ideal candidate for benefitting from curriculum learning. Moreover, curriculum learning has been shown to have an effect similar to unsupervised pre-training, acting as a form of regularization [6]. For theorem proving, curriculum learning allows LeanAgent to build a foundational understanding of mathematics before attempting more complex theorems, naturally leading to continuous improvement. This avoids suboptimal proof strategies early in training. Furthermore, this leads to more robust and generalizable proof techniques that work across a broader range of theorems, explaining LeanAgent’s sorry theorem proving performance. Moreover, the ability to act as a regularizer means curriculum learning prevents the model from overfitting to specific types of proofs or mathematical domains. This allows for continuous generalizability, explaining LeanAgent’s superior lifelong learning metric scores.

### v4

@@ A.6 Curriculum Learning Analysis
Curriculum Learning. Prior work suggests that curriculum learning guides the learner towards better local minima in non-convex optimization problems [6]. Crucially, theorem proving involves navigating a highly non-convex optimization landscape, especially for complex mathematical statements. The space of possible proofs is vast and complex, with many potential dead ends and suboptimal solutions to parts of a proof. This makes theorem proving an ideal candidate for benefitting from curriculum learning. Moreover, curriculum learning has been shown to have an effect similar to unsupervised pre-training, acting as a form of regularization [6]. For theorem proving, curriculum learning allows LeanAgent to build a foundational knowledge base of mathematics before attempting more complex theorems, naturally leading to continuous improvement. This avoids suboptimal proof strategies early in training. Furthermore, this leads to more robust and generalizable proof techniques that work across a broader range of theorems, explaining LeanAgent’s sorry theorem proving performance. Moreover, the ability to act as a regularizer means curriculum learning prevents the model from overfitting to specific types of proofs or mathematical domains. This allows for continuous generalizability, explaining LeanAgent’s superior lifelong learning metric scores.
@@
A.7 Theorem Proving Performance Score (TPPS) Analysis
In this section, we discuss the rationale behind the TPPS and acknowledge its limitations. The field of theorem proving, especially in a lifelong learning context, currently lacks standardized performance metrics. TPPS represents an initial attempt to address this gap, but it is important to recognize its preliminary nature and potential shortcomings.
The core idea behind TPPS is to assign a higher reward for newly proven theorems, aligning with the objectives of lifelong learning. However, this metric can be susceptible to artifacts that artificially inflate performance. For instance, several theorems in the PFR repository were proven on a technicality due to placeholder statements of “$0=1$”, which could then be used to prove other “sorry” theorems through the principle of ex falso quodlibet. This highlights a critical weakness in our current approach and underscores the need for more robust and well-tested benchmarks.
Furthermore, TPPS does not account for the difficulty of the theorems being proved. While we initially considered using proof length as a proxy for difficulty (e.g., $e^{S}$ where $S$ is the number of proof steps), we found this approach problematic. LeanAgent often generates shorter proofs for difficult theorems (e.g., the one-line proof for the wedderburn theorem compared to the 175-line human proof), making proof length an unreliable indicator of theorem difficulty.
These issues expose broader challenges in evaluating theorem-proving systems. The field lacks standardized benchmarks, necessitating carefully curated sets of theorems with varying difficulties across different mathematical domains. Developing reliable metrics to assess theorem difficulty beyond simple measures like proof length or statement complexity is crucial. Future benchmarks should prevent exploitation of technicalities. Moreover, metrics should consider the mathematical significance of proven theorems, not just their quantity. Finally, ensuring that performance metrics are meaningful and comparable across different mathematical repositories is essential for consistent evaluation.
In light of these considerations, we acknowledge that TPPS, while a step towards quantifying theorem-proving performance in a lifelong learning setting, has limitations. Future work should focus on developing more sophisticated and robust evaluation frameworks that address these challenges. By highlighting these issues, we hope to contribute to the ongoing discussion on how best to evaluate and compare theorem-proving systems, particularly in the context of lifelong learning. The development of standardized, reliable metrics will be crucial for measuring progress and guiding future research in this field.